OpenStudy (anonymous):

Solve sin(x) - √(1 - 3sin²(x)) = 0 given that 0° ≤ x ≤ 360°.

OpenStudy (irishboy123):

\[\sin x - \sqrt{1 - 3 \sin^2 x} = 0\] \[\sin x =\sqrt{1 - 3 \sin^2 x}\] does that get the ball rolling?

OpenStudy (anonymous):

I did that and then squared both sides

OpenStudy (irishboy123):

cool. so problem solved? or do you need more?

OpenStudy (er.mohd.amir):

squaring both side sin^2 x=1-3 sin^2 x 4 sin^2 x=1 sin x=+1/2 or -1/2 x=30 degree or 210 degree

OpenStudy (irishboy123):

good job